# Explainability of Large Language Models using SMILE: Statistical Model-agnostic Interpretability with Local Explanations

### v1

@@ Abstract
Large language models like GPT, LLAMA, and Claude have become incredibly powerful at generating text, but they are still black boxes, so it is hard to understand how they decide what to say. That lack of transparency can be problematic, especially in fields where trust and accountability matter. To help with this, we introduce SMILE, a new method that explains how these models respond to different parts of a prompt. SMILE is model-agnostic and works by slightly changing the input, measuring how the output changes, and then highlighting which words had the most impact. Create simple visual heat maps showing which parts of a prompt matter the most. We tested SMILE on several leading LLMs and used metrics such as accuracy, consistency, stability, and fidelity to show that it gives clear and reliable explanations. By making these models easier to understand, SMILE brings us one step closer to making AI more transparent and trustworthy.
@@ 1 Introduction
Large Language Models (LLMs) have become powerful tools in recent years. They can perform tasks ranging from answering questions with remarkable accuracy [1] to drafting essays and other written content with coherence and style [2], creating art that spans various mediums and styles [3], coding across multiple programming languages [4], debugging software [4], writing technical documentation [5], and even generating algorithms [1].
LLMs function by leveraging deep learning architectures, particularly the Transformer model. This model uses self-attention mechanisms to process and generate text by analysing relationships between words in large data sets [6]. These models are trained in various corpora, including books, articles, code, and web content, allowing them to learn language patterns, context, and syntax on an unheard scale [7].
Their capabilities continue to grow, turning them into essential tools for a wide range of tasks [2]. These models, trained in various texts, have transformed how we interact with technology and how machines understand and generate human language [6]. LLMs such as GPT, BERT, and Claude leverage the Transformer architecture to process and generate language based on patterns learned from extensive datasets [7]. Although they excel in many NLP tasks, understanding their strengths and limitations is the key to using them effectively and responsibly [8].
These models continue to evolve, pushing the boundaries of AI applications in language understanding and generation [2]. However, the nature of these models is known as black boxes and lacks transparency [9]. One of their important problems is how these models decide which words or parts of input to focus on, which remains a complex and fascinating process [10].
At their core, LLMs rely on patterns and relationships learned during training to decide which input elements are the most important [6]. This process is more than just identifying individual keywords; it is about picking up sophisticated connections and context that shape the model’s responses [10]. Understanding how this works is crucial: it explains why these models excel in some scenarios but struggle in others. It also highlights why they can occasionally produce responses that are biased, irrelevant, or even harmful [11].
By studying how LLMs emphasise words and phrases, we can better understand how they work [10]. This can help us make them more transparent and reliable, ensuring they focus on what truly matters in a given context. It also allows us to address potential problems, such as biases or alignment problems, making these tools safer and more ethical [11].
As LLMs continue to integrate into our daily lives and decision-making processes, understanding their inner workings is more important than ever [9]. Research on how they prioritise language helps us refine their performance, ensure ethical use, and unlock their full potential to benefit society [11] responsibly.
To solve this problem, we propose a novel approach of SMILE (Statistical Model-agnostic Interpretability with Local Explanations) to extract a visual heat map that shows the influence of each word in responding to these models [12]. This method extends the foundation laid by LIME [13] by incorporating Empirical Cumulative Distribution Function (ECDF) statistical distances, significantly improving robustness and making interpretability more effective.
SMILE produces visual heatmaps that reveal the impact of individual words or phrases in text input on the model’s output as illustrated in Fig 1. For example, when a language model processes an instruction or query, the heatmap highlights the importance of specific terms in influencing the outcome. This offers users a clear visual representation of how the model interprets and prioritises different parts of the input text [10].
To ensure a thorough evaluation of our approach, we employ a suite of metrics, including accuracy, fidelity, stability, and consistency [14]. These metrics demonstrate that SMILE provides reliable insights into model behaviour [15].
The contributions of our work include the following.
  • Introducing a New Interpretability Framework: SMILE is a cutting-edge method that visualises the role of individual words in shaping model output, offering a deeper understanding of how LLMs process language [12].
  • Promoting Transparency and Trust in LLMs: By visually linking text input to model decisions, SMILE enhances transparency, making LLMs more trustworthy and enabling their practical use in high-stakes domains such as healthcare, education, and legal systems.
  • Comprehensive Evaluation for Robustness: Our evaluation leverages metrics such as stability, fidelity, accuracy, and consistency, establishing a new standard for evaluating the interpretability and reliability of language models [16].
@@ 2 Literature review
This section summarises the most recent advances in large language models (LLMs) and their growth in various applications. We also discuss Explainable Artificial Intelligence (XAI), a field dedicated to making AI systems more transparent and easier to understand. Finally, we explore how these two areas come together, focusing on the ongoing efforts to create explainable large language models. We discuss the key challenges involved and the techniques being developed to improve how we interpret and trust these powerful models.
@@ 2.1 Large Language Models
Large language models (LLMs) have become increasingly common thanks to significant strides in deep learning, better access to data, and more powerful computing resources. These models are trained on vast amounts of text and can generate responses that sound natural, answer questions, and support a wide range of language-based tasks. Over time, several notable LLMs have helped push the field forward, each bringing new ideas and improvements.
GPT (OpenAI) [1] is one of the most well-known autoregressive transformer-based models, capable of generating coherent and contextually rich text. Google’s T5 [17] introduced the text-to-text framework, where all NLP tasks are framed as text generation problems, improving transfer learning and efficiency. Similarly, PaLM (Google) [18] leverages the Pathways architecture for large-scale language modelling, optimising multitask learning across different NLP tasks.
Anthropic’s Claude [19] focuses on ethical AI and safe interactions designed to minimise harmful outputs while maintaining interpretability. Google’s Gemma 2 [20] is optimised for logical reasoning and structured data interpretation, making it a strong candidate for academic and enterprise use cases. Meta’s LLaMA [5] is an efficient, open-source model tailored for various NLP applications, offering researchers an accessible yet powerful alternative to proprietary models.
EleutherAI’s GPT-NeoX [21] is an open-source alternative to GPT, providing robust text generation capabilities while allowing customisation. AI21 Labs’ Jurassic-2 [22] is designed for large-scale enterprise applications, emphasising contextual understanding and user intent. Cohere’s Command R+ [23] integrates retrieval-augmented generation, improving fact-based responses and minimising hallucinations.
Furthermore, models like BERT (Google) [24] have NLP through bidirectional context understanding, significantly improving sentiment analysis and question-answering tasks. Microsoft’s Phi-4 focuses on efficiency and cost-effectiveness, delivering competitive performance with fewer parameters. Mistral and Alibaba’s Qwen 2.5 push the boundaries of model adaptability and fine-tuning capabilities, making them suitable for multilingual and domain-specific applications.
DeepSeek-LLM [25] is another promising open-source model, competing with top-tier models regarding benchmark performance. Models such as Mixtral and StableLM have been developed with efficiency and specific NLP task performance in mind, ensuring a balance between accuracy and computational demands.
Recent developments have introduced models such as DeepSeek-R1, which focused on advanced reasoning capabilities, and Qwen2.5-Max, which explores large-scale Mixture-of-Experts (MoE) models to enhance efficiency and performance. OpenAI’s o3-mini pushes the frontier of cost-effective reasoning, balancing performance with computational efficiency. DeepSeek-V3, the first open-sourced GPT-4o level model, represents a significant step forward in open-source LLM development.
@@
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications:
@@
Category Models Description General-Purpose GPT-4 [26], GPT-4o [27], Claude 3 [28], Gemini 1.5 [29], Command R+ [30], Mistral [31] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [32], Mistral [31], Mixtral [33], Falcon 180B [34], DeepSeek-V3 [35], OpenChat [36] Customizable models for research applications Retrieval-Augmented Command R+ [30], GPT-4 Bing [26], Claude 3 RAG [28], Gemini Search [29] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [28], Gemini 1.5 [29], GPT-4o [27], Grok-1.5 [37] Specialized in processing long-form text Advanced Reasoning GPT-4 [26], Claude 3 Opus [28], Gemini 1.5 [29], Mixtral [33], Grok-1.5 [37] Logical reasoning, multi-task learning, problem-solving Ethically-Aligned Claude 3 [28], Gemini [29], GPT-4 [26], Grok [37] Emphasize safety, reducing bias, ethical AI Mixture-of-Experts (MoE) Mixtral [33], Qwen2.5-Max [38], GShard [39], Switch Transformer [40] Improved efficiency and performance via MoE Multimodal GPT-4o [27], Gemini 1.5 [29], Claude 3 [28], Kosmos-2 [41], Grok-1.5V [37] Capable of processing both text and visual inputs Multilingual BLOOM [42], XGLM [43], mBERT [44], XLM-R [45], ByT5 [46] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [47], Galactica [48], Legal-BERT [49], StarCoder2 [50], Code LLaMA [51] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [52], Vicuna [53], ChatGLM [54], Baize [55], OpenChat [36] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [56], TinyLLaMA [57], DistilGPT2 [58], GPT-2 [59], LLaMA 2 7B [60] Designed for fast inference and deployment on edge devices
@@ 2.2 Explainable AI (XAI)
The increasing complexity of machine learning models has made understanding how they make decisions more important than ever. Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [61, 14].
Explainability methods can be broadly categorised as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [9].
@@
Global explanations provide an overarching view of the model’s decision-making patterns across an entire dataset, helping to identify trends and biases [62, 64]. In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs. LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [13].
@@
LIME has become one of the most widely used explainability techniques, but over time, researchers have introduced various improvements to address its limitations and expand its usefulness across different applications. As shown in Figure 3, enhancements to LIME can be grouped into four main categories:
@@
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [65], S-LIME [66], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behaviour, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimisation strategies to make LIME more efficient and stable. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
@@
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [67], TS-MULE for time series data [68], Graph LIME for graph-structured data [69], Sound LIME for audio data [70], and B-LIME for ECG signal data [71] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to improve LIME’s approach, including using alternative surrogate models, fine-tuning distance parameters, optimising sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
Figure 3: Key strategies for enhancing LIME, including improvements in sampling, distance parameters, surrogate models, and optimisation.

### v2

# Explaining Large Language Models with gSMILE: Generative Statistical Model-Agnostic Interpretability with Local Explanations

@@ Abstract
Large Language Models (LLMs) such as GPT, LLaMA, and Claude have achieved remarkable performance in text generation, yet their decision-making processes remain opaque. This lack of transparency poses challenges in domains where trust and accountability are critical. In this work, we introduce gSMILE (Generative SMILE), a generalized and LLM-oriented extension of the previously proposed SMILE framework, for interpreting how these models respond to specific parts of an input prompt.
gSMILE is model-agnostic and operates by introducing controlled perturbations to the input, observing the resulting output changes, and identifying which input words have the most significant impact on the output. This process enables the generation of intuitive heatmaps that visually highlight influential words in the prompt. We evaluate gSMILE across several leading large language models (LLMs) using metrics such as accuracy, consistency, stability, and fidelity, demonstrating its ability to provide precise, robust, and reliable explanations. By making LLMs more interpretable, gSMILE takes an important step toward building more transparent and trustworthy AI systems.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT, Claude, and LLaMA, have significantly advanced the capabilities of natural language processing systems [1, 2, 3]. These models, based on the Transformer architecture [4], are trained on extensive corpora and exhibit remarkable performance across a wide range of tasks, including text generation, summarisation, code completion, and dialogue reasoning [5]. However, despite their widespread adoption, LLMs remain largely opaque in how they make predictions, leading to concerns about transparency, trust, and safety, especially in high-stakes domains such as healthcare and law [6, 7].
The black-box nature of LLMs limits our ability to understand which parts of an input prompt have the most significant influence on the model’s output. This lack of interpretability hinders debugging, ethical assessment, and responsible deployment [8, 9]. Interpretability methods, especially those that operate post hoc and in a model-agnostic fashion, offer a way to open these black boxes by explaining the model’s behavior without requiring access to internal parameters [10].
In this work, we propose gSMILE, a generalized and LLM-oriented extension of the SMILE (Statistical Model-Agnostic Interpretability with Local Explanations) framework [11]. gSMILE interprets model outputs by perturbing input prompts, measuring output shifts using Wasserstein distance, and fitting a local linear surrogate model. The result is a weighted heatmap that visualizes how individual words in a prompt contribute to the model’s response.
We evaluate gSMILE across multiple large language models (LLMs) and prompt types using quantitative metrics, including fidelity, consistency, stability, and accuracy. Our findings demonstrate that gSMILE yields stable, interpretable, and reliable attributions, rendering it suitable for practical applications in real-world scenarios involving instruction-based generation.
The key contributions of this paper are:
  • Token-level interpretability for LLMs: We present gSMILE, a model-agnostic method that quantifies the influence of each input token on the output of large language models using perturbation-based analysis and Wasserstein distance.
  • Visual attribution through weighted heatmaps: The framework generates intuitive heatmaps that highlight high-impact tokens, enabling human-aligned interpretation of model behavior.
  • Robust cross-model evaluation: We evaluate gSMILE on multiple instruction-tuned LLMs using diverse metrics, accuracy, fidelity, stability, and consistency that demonstrate its reliability and adaptability.
@@ 2 Literature review
This section provides an overview of recent advancements in large language models (LLMs) and their increasing impact across various applications. We also introduce Explainable Artificial Intelligence (XAI), which aims to make AI models more transparent, interpretable, and trustworthy. As LLMs grow in complexity and usage, the need to understand their decision-making processes becomes more critical. Therefore, we highlight the emerging efforts to develop explainable large language models (LLMs), outlining key challenges and recent techniques that aim to improve their interpretability and foster greater trust in their outputs.
@@ 2.1 Large Language Models
Large language models (LLMs) have become increasingly prevalent due to significant advancements in deep learning [12], improved access to large-scale datasets, and the growing availability of powerful computing resources [13]. These models are typically trained on massive corpora of text. They are capable of generating fluent, contextually appropriate responses across a wide variety of language-based tasks, including question answering, summarisation, translation, and content generation. Over time, several influential large language models (LLMs) have helped push the field forward, each introducing novel architectural innovations [4], training strategies [14], or alignment techniques [15] that have contributed to their growing capabilities and broader adoption.
Among the most influential models, GPT (OpenAI) [1] stands out as a leading autoregressive transformer-based architecture capable of generating coherent and contextually rich text. It set a new standard for generative language models and introduced the broader public to powerful applications of AI through tools like ChatGPT. Expanding on this paradigm, Google’s T5 [14] reframed all NLP tasks as text-to-text problems, streamlining the transfer of learning across diverse tasks and contributing to more unified and efficient architectures. Palm (Google) [16], with its Pathways design, extended the multitask learning capability at scale, allowing a single model to generalize effectively across different NLP applications.
In parallel, Anthropic’s Claude [17] introduced a new approach centered on safe and aligned AI behavior. Claude utilizes constitutional AI principles to ensure its outputs are consistent with predefined ethical standards, thereby promoting greater transparency and reducing the risk of harm. More recently, Google’s Gemma 2 [18] has focused on logical reasoning and structured data interpretation, offering a competitive option for research-driven and enterprise-level deployment. Meta’s LLaMA [19] series continues to draw attention as a scalable, open-source alternative, providing the research community with accessible tools that deliver high performance across a range of benchmarks.
Community-driven efforts have also enriched the landscape. EleutherAI’s GPT-NeoX [20] was developed to provide an open-source alternative to proprietary GPT models, offering researchers greater flexibility in experimentation and deployment. AI21 Labs’ Jurassic-2 [21] targets commercial-scale applications with a focus on contextual awareness and enterprise integration. Meanwhile, Cohere’s Command R+ [22] integrates retrieval-augmented generation (RAG), enhancing factual consistency and minimizing hallucinations in the generative output — a key concern for the practical adoption of large language models (LLMs).
Additionally, transformer-based models, such as BERT [23], have revolutionized NLP through bidirectional encoding, providing substantial improvements in tasks like question answering, named entity recognition, and sentiment analysis. Efficiency-focused models, such as Microsoft’s Phi-4, highlight a growing trend toward parameter-efficient design without sacrificing core performance, making large language models (LLMs) more accessible for low-resource environments. Mistral and Qwen 2.5 (by Alibaba) continue this direction by offering fine-tuning-friendly models with strong multilingual and domain adaptation capabilities.
Several recent open-source contenders further enrich the ecosystem. DeepSeek-LLM [24] delivers performance comparable to proprietary models while remaining open and extensible. Mixtral and StableLM focus on task efficiency and targeted performance, ensuring their utility in production environments. DeepSeek-R1 has been noted for its emphasis on reasoning, reflecting a broader shift from general language generation to models optimized for critical thinking, chain-of-thought prompting, and interpretability.
Cutting-edge advancements have introduced even more specialized architectures. Qwen2.5-Max utilizes large-scale Mixture-of-Experts (MoE) models to enhance throughput and scalability. OpenAI’s o3-mini emphasizes low-latency, cost-effective deployment while retaining high-quality output — ideal for mobile and edge computing. Perhaps most notably, DeepSeek-V3 represents a significant milestone by achieving GPT-4o-level performance in an open-source setting, suggesting that open-source models may soon rival or surpass closed models in both capability and adoption.
@@
The following models in Fig. 2 are explicitly designed for text-to-text tasks, meaning they take text input and generate text output across various NLP applications. These models continue to define the state of the art in machine-generated language, with a growing focus on transparency, efficiency, and alignment with human values.
@@
Category Models Description General-Purpose GPT-4 [25], GPT-4o [26], Claude 3 [27], Gemini 1.5 [28], Command R+ [29], Mistral [30] Versatile for wide range of NLP tasks Open-Source LLaMA 3 [31], Mistral [30], Mixtral [32], Falcon 180B [33], DeepSeek-V3 [34], OpenChat [35] Customizable models for research applications Retrieval-Augmented Command R+ [29], GPT-4 Bing [25], Claude 3 RAG [27], Gemini Search [28] Enhance information retrieval accuracy, reduce misinformation Long-Text Processing Claude 3 [27], Gemini 1.5 [28], GPT-4o [26], Grok-1.5 [36] Specialized in processing long-form text Advanced Reasoning GPT-4 [25], Claude 3 Opus [27], Gemini 1.5 [28], Mixtral [32], Grok-1.5 [36] Logical reasoning, multi-task learning, problem-solving Safety-Focused Design Claude 3 [27], Gemini [28], GPT-4 [25], Grok [36] Emphasize alignment with safety rules and bias mitigation Mixture-of-Experts (MoE) Mixtral [32], Qwen2.5-Max [37], GShard [38], Switch Transformer [39] Improved efficiency and performance via MoE Multimodal GPT-4o [26], Gemini 1.5 [28], Claude 3 [27], Kosmos-2 [40], Grok-1.5V [36] Capable of processing both text and visual inputs Multilingual BLOOM [41], XGLM [42], mBERT [43], XLM-R [44], ByT5 [45] Optimized for cross-lingual understanding and translation Domain-Specific Med-PaLM 2 [46], Galactica [47], Legal-BERT [48], StarCoder2 [49], Code LLaMA [50] Tailored for specific tasks such as medical, legal, or code Instruction-Tuned Alpaca [51], Vicuna [52], ChatGLM [53], Baize [54], OpenChat [35] Tuned for better instruction-following and chat capabilities Lightweight / Edge Phi-2 [55], TinyLLaMA [56], DistilGPT2 [57], GPT-2 [58], LLaMA 2 7B [59] Designed for fast inference and deployment on edge devices
@@ 2.2 Explainable AI (XAI)
The increasing complexity of machine learning models has made understanding how they make decisions more critical than ever. Ensuring transparency, fairness, and reliability is crucial, so explainability has become a key focus in AI research [60, 61].
Explainability methods can be broadly categorized as intrinsic or post-hoc. Intrinsic methods involve inherently interpretable models due to their simplicity, such as linear regression or decision trees [62]. These models are often preferred in applications where transparency is critical, but they may lack the predictive power of more complex models [6].
@@
Global explanations offer an overarching view of the model’s decision-making patterns across an entire dataset, facilitating the identification of trends and biases [62, 66]. In contrast, local explanations focus on specific predictions, providing a detailed breakdown of how individual inputs influence model outputs. LIME and its numerous enhancements fall under local explainability methods and are designed to provide granular, case-specific interpretations [63].
@@
LIME has become one of the most widely used explainability techniques; however, over time, researchers have introduced various improvements to address its limitations and expand its applicability across different applications. As shown in Figure 3, enhancements to LIME can be grouped into four main categories.
@@
A key improvement in LIME’s explanations is the refinement of distance measures used to weigh data points. For instance, SMILE uses statistical techniques to generate more consistent and reliable explanations. However, while these improvements enhance reliability, they also come with increased computational complexity. Another approach to improving LIME focuses on upgrading its simple linear models with more sophisticated alternatives. Methods such as Q-LIME [67], S-LIME [68], Bay-LIME, and ALIME use more advanced surrogate models to provide explanations that better reflect the underlying model’s behavior, capturing non-linear relationships that the original method might miss. Furthermore, researchers have explored optimization strategies to enhance the efficiency and stability of LIME. Techniques like OptiLIME and G-LIME work to strike a balance between accuracy and computational cost, ensuring that explanations remain reliable while keeping the process computationally manageable.
@@
Recent research has extended LIME to support specific data types and improve its interpretability. Adaptations such as PointNet LIME for 3D point cloud data [69], TS-MULE for time series data [70], Graph LIME for graph-structured data [71], Sound LIME for audio data [72], and B-LIME for ECG signal data [73] demonstrate how LIME can be tailored to diverse applications. In addition, several modifications have been proposed to enhance LIME’s approach, including the use of alternative surrogate models, fine-tuning distance parameters, optimizing sampling techniques, and refining computational efficiency to improve both interpretability and accuracy.
Other extensions include DSEG-LIME [74], which tailors LIME for segmentation models by generating perturbations at the object level rather than the pixel level, improving interpretability in image segmentation tasks. SLICE [75] introduces structured sampling and clustering to ensure diverse and representative local explanations, reducing redundancy in LIME samples. Stratified LIME [76] improves fidelity by stratifying the input space and generating more targeted local samples for different subpopulations. US-LIME [77] employs uncertainty sampling to guide instance selection, increasing the reliability and faithfulness of local explanations. SS-LIME [78] uses self-supervised learning to identify important regions in the feature space and provide richer explanations. K-LIME [79] incorporates kernel-based methods into LIME’s regression step to model more complex local decision boundaries. Attention-LIME [80] leverages attention mechanisms to prioritize features during the explanation, aligning more closely with deep learning models. BMB-LIME [81] employs a boosted multi-branch approach to combine multiple local models, thereby enhancing consistency and robustness.
These methods aim to enhance interpretability across various domains, including segmentation, survival analysis, multimodal learning, and attention-based contexts. Additionally, a wide range of other extensions have been developed to enhance LIME’s capabilities across domains and explanation strategies, including: LSLIME, LIME-Aleph, Anchors [82], LIME-SUP, LORE, ILIME, Aleph LIME, bLIMEy, NormLIME, DLIME, LIMEtree, MPS-LIME, Explain Explore, survLIME, survLIME-inf, CBR-LIME, GuidedLIME, audioLIME, MeLIME, C-LIME, gLIME, DIME, s-LIME, LIMESegment, UnRAvEL-LIME, Kernel-LIME, Specific-Input LIME, ExpLIMEable, GLIME, GMM-LIME, SEGAL, and Sig-LIME, reflecting the diverse and rapidly evolving landscape of post-hoc explainability [83].
Figure 3: Taxonomy of LIME enhancement strategies: categorized into sampling techniques, surrogate model improvements, distance parameter adjustments, and optimization approaches.